OpenStudy (anonymous):

The x-intercepts of the graph of ƒ(x) = x(x + 1)(x – 5)(x – 7) are ,

OpenStudy (campbell_st):

well to find the zeros solve the factors to find the values of x that make each factor zero. so solve x = 0 x + 1 = 0 x - 5 = 0 x -7 = 0 hope it helps

OpenStudy (anonymous):

It actually does help a lot!! Thank you so much!